Your company has three million shares of common stock outstanding with a current market price of $28. RE = 14.5%. There are also 50,000 bonds outstanding with an 8 percent annual coupon, 18 years to maturity, and a current price of $940 per bond. What is the WACC for your company assuming a tax rate of 35%. 16
Answer #1

Value of shares =3000000*28 =84000000

Value of debt =50000*940 =47000000

Total of shares and debt =131,000,000

After-tax cost of debt of bonds:

=RATE(18,8%*1000,-940,1000)*(1-0.35)

=5.64%

WACC:

=14.5%*(84/131)+5.64%*(47/131)

=11.32%